OpenStudy (anonymous):

If 10^x=3, find the value of 10^(2x+1).

OpenStudy (anonymous):

log(3)=x, so 10^(2log(3)+1) would be your answer you can further simplify to 10^(log(9)+1)

OpenStudy (anonymous):

Can't it be solved without using log?

OpenStudy (anonymous):

Switching from exponential to logarithmic is the easiest way to find solutions to these types of problems.

OpenStudy (anonymous):

But we are required to ude exponential method.

OpenStudy (anonymous):

10^(2x+1) = 10^(2x) * 10^(1) = (10^x)^2 * 10.

OpenStudy (anonymous):

*use

OpenStudy (anonymous):

Please mark your question as closed once it is solved

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

love the penguins by the way

OpenStudy (anonymous):

Thank you! ^_^